OpenStudy (anonymous):

Please help me with this! Simplify the following quotient of complex numbers into the form a+bi. -9 + 2i / -8 - 9i

OpenStudy (mertsj):

In this type of problem, always multiply by the conjugate of the denominator.

OpenStudy (anonymous):

unfortunately... you just spoke a completely different language to me

OpenStudy (mertsj):

We are trying to set up a situation in the denominator which will eliminate the radical. If we can have a (a-b)(a+b) product, that will do the trick.

OpenStudy (mertsj):

Since the denominator is -8-9i we need to multiply that by -8+9i which is called the conjugate of -8-9i

OpenStudy (anonymous):

so -9+2i / -8-9i(-8+9i)?

OpenStudy (mertsj):

Must multiply the numerator by the same thing.

OpenStudy (mertsj):

\[\frac{-9+2i}{-8-9i}\times\frac{-8+9i}{-8+9i}\]

OpenStudy (anonymous):

-9+2i(-8+9i) / -8-9i(-8+9i)

OpenStudy (mertsj):

No. The first denominator is -8-9i

OpenStudy (anonymous):

is that not what I put?

OpenStudy (mertsj):

Just look at you next to last post.

OpenStudy (anonymous):

ugh.... Im so bad at math

OpenStudy (anonymous):

i got -72+18i

OpenStudy (mertsj):